OpenStudy (anonymous):

HELP HELP HELP PLEASE? Can someone explain or tell me how to do this math question?? You invest $4,100 in an account that pays an interest rate of 7.25%, compounded continuously. Calculate the balance of your account after 8 years. Round your answer the nearest hundredth

OpenStudy (gabebae):

Use the interest rate formula \[I = Ptr\]

OpenStudy (anonymous):

is that like A=Pe^(rt) @GabeBae

OpenStudy (anonymous):

@Hoslos

OpenStudy (gabebae):

yes now you just plug the numbers in that you have

OpenStudy (anonymous):

@GabeBae , the PTR you gave is different from \[P ^{tr}\]

OpenStudy (anonymous):

crooect that is the one that i meant to put up @Hoslos

OpenStudy (anonymous):

correct*

OpenStudy (gabebae):

sorry

OpenStudy (anonymous):

so i plugged them in and i got 7322.75 but it seem wrong to me /: @Hoslos

OpenStudy (anonymous):

Ok. Let me work it out.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

I realized that year after year, the values are\[4100\] \[4100+(4100*0.0725)=4397.25\] \[4397.25+(0.0725*4397.25)=4716...\] And so on. Conclusively, the formula for compound interest is:\[Total=PrincipalAmount*(1+interest)^t=4100*(1+0.0725)^8=7200\] to the nearest hundreth.

OpenStudy (anonymous):

oh okay i see now @Hoslos

OpenStudy (anonymous):

Good.

OpenStudy (anonymous):

yes thank you @Hoslos